An electromagnetic wave in free space has an electric- eld vector E = f(t - z/c0).x, where ^x is a unit

vector in the x direction, and f(t) = e^-t^2=2

e^j2pv0t, where y is a constant. Describe the physical nature

of this wave and determine an expression for the magnetic- eld vector.


1
Expert's answer
2021-06-28T17:02:02-0400

Gives

"E=f(t-\frac{z}{c_0})\hat{x}"

"f(t)=e^{-t^2},y=2e^{j2\pi v_0t}=2e^{jwt}"

We know that wave equation

"y=Ae^{(jw_0t+\phi)}"

Where A=amplitude

"w_0=" Anguler frequency

"\phi" =phase

"y=2e^{j2\pi v_0t}=2e^{jwt}\rightarrow(1)"

equation (1) is showing wave equation

equation (1)is

Phase

"\phi" =0°

Amplitude=2

Anguler frequency

"w=2\pi v_0"

Magnetic field vector

"B=\frac{E}{c}"

"B=\frac{E}{c}=\frac{f(t-\frac{z}{c_0})\hat{x}}{c}"

Where

"f(t)=e^{-t^2}"